OpenStudy (anonymous):

Which of the following best defines 3^2/3? (New Question at the Bottom) Square root of 9 Cube root of 9 Cube root of 3 Square root of 3

OpenStudy (anonymous):

@pooja195

OpenStudy (anonymous):

\[\huge 3^{2/3} \implies \sqrt[3]{3^2}\]

OpenStudy (anonymous):

The general rule \[\huge x^{\frac{ m }{ n }} =\sqrt[n]{x^m}\]

OpenStudy (anonymous):

Right, then what?

OpenStudy (anonymous):

What does \[\sqrt[3]{3^2 }\] mean?

OpenStudy (anonymous):

Or lets do this step by step what is \[\large 3^2\]

OpenStudy (anonymous):

Uh, do you multiply 3 x 2?

OpenStudy (anonymous):

Oh my bad. It's 9

OpenStudy (anonymous):

No, these are exponents the number in the exponent implies how many times you're multiplying it by it self, so \[x^1 = x\]\[x^2 = x \times x\]\[x^3 = x \times x \times x\]

OpenStudy (anonymous):

Right!

OpenStudy (anonymous):

So now we have \[\large \sqrt[3]{9}\]

OpenStudy (anonymous):

Note a square root is \[\huge \sqrt{x}\] so what does \[\huge \sqrt[3]{x}\] mean

OpenStudy (anonymous):

\[\sqrt{x} = x^{1/2}~~~~~~\sqrt[3]{x} = x^{1/3}\]

OpenStudy (anonymous):

I'm not really sure,

OpenStudy (anonymous):

Well a square root is just \[\huge \sqrt{x} = x^{1/2}\] and a quad root is \[\huge \sqrt[4]{x} = x^{1/4}\] there's a pattern going on think you notice it yet?

OpenStudy (anonymous):

\[\huge \sqrt[3]{x} = x^{1/3}\] what is this called?

OpenStudy (anonymous):

Look at your options if you have to

OpenStudy (anonymous):

Is it a cube root?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

So that leaves me to either B or C

OpenStudy (anonymous):

Think about it, we already did the first part so we have \[\huge \sqrt[3]{9}\]

OpenStudy (anonymous):

Emm, is it B?

OpenStudy (anonymous):

Sounds good..

OpenStudy (anonymous):

Cool thanks! You think you can help me with another?

OpenStudy (anonymous):

Part A: Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days. Part B: Write the equation obtained in Part A using function notation. Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals.

OpenStudy (anonymous):

... Or I can just close this question and open another...